OpenStudy (anonymous):

Kelly loves to eat chocolate pizza. Each square inch of chocolate is 30 calories. She orders a slice of chocolate pizza from a pizza with a diameter of 12 inches that is cut with a central angle of 40 degrees. How many calories does she eat? about 3,393 calories about 113 calories about 13 calories about 377 calories

OpenStudy (bahrom7893):

Yum chocolate pizza!

OpenStudy (callisto):

area of the pizza she eats = (pi) (12/2)^2 x 40/360 = 4 pi in^2 calories she eats = 4 (pi) x 30 = 377

OpenStudy (anonymous):

Haha!

OpenStudy (anonymous):

are you using arc length or arc area formula?

OpenStudy (callisto):

area of the sector \[=\pi r ^{2} \theta/360\]

OpenStudy (anonymous):

Gotcha! I forgot to multiply by the 30. I originally get the 4. Thanks!

OpenStudy (callisto):

welcome :)